Josh Company had 200,000 ordinary shares outstanding on December 31, 2017. The company issued 20,000 shares on April 1, 2018 and retired 10,000 shares on September 1, 2018. Josh company recorded net income for the year ended December 31, 2018 of $300,000 after a loss on discontinued operations of $35,000 (net of tax).

Calculate EPS (Earnings per share)
Answer #1

Solution:

Weighted average outstanding number of shares = 200000*12/12 + 20000*9/12 - 10000*4/12 = 200000 + 15000 - 3333.33333

= 211,666.6667

EPS = Net income / Weighted average outstanding number of shares = $300000 / 211666.6667 = $1.42 (rounded to 2 decimals)